# AdvI2I: Adversarial Image Attack on Image-to-Image Diffusion models

### v2

@@ Abstract
Recent advances in diffusion models have significantly enhanced the quality of image synthesis, yet they have also introduced serious safety concerns, particularly the generation of Not Safe for Work (NSFW) content. Previous research has demonstrated that adversarial prompts can be used to generate NSFW content. However, such adversarial text prompts are often easily detectable by text-based filters, limiting their efficacy. In this paper, we expose a previously overlooked vulnerability: adversarial image attacks targeting Image-to-Image (I2I) diffusion models. We propose AdvI2I, a novel framework that manipulates input images to induce diffusion models to generate NSFW content. By optimizing a generator to craft adversarial images, AdvI2I circumvents existing defense mechanisms, such as Safe Latent Diffusion (SLD), without altering the text prompts. Furthermore, we introduce AdvI2I-Adaptive, an enhanced version that adapts to potential countermeasures and minimizes the resemblance between adversarial images and NSFW concept embeddings, making the attack more resilient against defenses. Through extensive experiments, we demonstrate that both AdvI2I and AdvI2I-Adaptive can effectively bypass current safeguards, highlighting the urgent need for stronger security measures to address the misuse of I2I diffusion models. The code is available at https://github.com/Spinozaaa/AdvI2I.
CAUTION: This paper includes sexually explicit imagery and discussions of pornography that may be disturbing or offensive to some readers.
@@ 1 Introduction
Recently, diffusion models have made significant strides in the domain of image synthesis, demonstrating their ability to produce high-quality images (Rombach et al., 2022; Zhang et al., 2023). However, these advancements have also raised significant ethical and safety concerns. Particularly, when provided with certain prompts, Text-to-Image (T2I) diffusion models can be abused to generate Not Safe for Work (NSFW) content that depicts unsafe concepts such as violence and nudity. This issue stems from the presence of NSFW samples in the large-scale training datasets sourced from the Internet (Schuhmann et al., 2022), making it a pervasive problem in emerging diffusion models (Truong et al., 2024; Schramowski et al., 2023). Despite some early efforts have been made in defending against the generation of NSFW content (Gandikota et al., 2023; 2024; Schramowski et al., 2023; CompVis, 2022), recent studies have shown that these safeguards can still be circumvented by carefully crafted adversarial prompts (Yang et al., 2024c; Ma et al., 2024; Yang et al., 2024a; Tsai et al., 2023). As a result, malicious users can exploit these models to generate NSFW images for unethical purposes.
While adversarial prompts present a notable risk to the generation safety of diffusion models, their Achilles’ heel lies in that such attacks work by changing the input text prompt, which can exhibit easily detectable patterns that distinguish them from natural prompts. Specifically, we applied four types of simple filters (perplexity filter, keyword filter, embedding filter and large language model (LLM) filter) to a range of adversarial prompt attacks (Zhuang et al., 2023; Kou et al., 2023; Tsai et al., 2023; Ma et al., 2024; Yang et al., 2024c), and found that even the simplest filters can effectively identify adversarial prompts from normal ones in most cases (see more detailed in Section 3.1). Notably, a naive perplexity filter can (on average) reduce the attack success rate (ASR) of adversarial prompts by , while using an LLM as the safety filter can reduce the ASR to under .
This suggests that adversarial text prompts can be identified, which means that diffusion models can reject generating images with such queries if detected. However, the new question is:
@@
In this work, we provide a negative answer to this question. We reveal the risk of adversarial images that can also induce diffusion models to generate NSFW images, which has not been well explored in previous research. We propose a framework named AdvI2I to demonstrate the effectiveness of such an attack on the Image-to-image (I2I) diffusion model, alerting the community to adversarial attacks from not only the prompt but also the image condition side. In addition to text prompts, I2I diffusion models conventionally utilize an image as a conditioning input. By leveraging adversarial images, attackers can induce the diffusion model to generate NSFW images. For example, an image of the president can be manipulated to depict nudity. Moreover, this method can bypass current defense mechanisms on diffusion models and thereby represents a significant but underexplored security vulnerability in this domain.
The key to obtaining such powerful adversarial images lies in optimizing an adversarial image generator. The optimization target is the denoised latent feature in the diffusion process. Given that the feature is influenced by both the image and text conditions, AdvI2I can encode the NSFW concept from the text embedding space to the parameter space of the adversarial images generator, enabling it to guide the model in generating NSFW content. Additionally, to further explore the efficacy of such adversarial attack under potential defenses, we propose a modified attack approach named AdvI2I-Adaptive. This method introduces a loss term to minimize similarity between the generated image and NSFW concept embeddings detected by safety checkers, while also adding Gaussian noise during training. By incorporating these adaptive elements, AdvI2I-Adaptive enhances the robustness of adversarial attacks against current defense measures, significantly amplifying the threat posed by adversarial images in I2I diffusion models. Our contributions are summarized as follows.
@@
By highlighting the risk of adversarial attacks from image conditions, this work raises awareness within the research community about the potential dangers of such attacks on diffusion models, urging further investigation and development of robust defenses.
2 Related Work
Adversarial Attack and Defense in T2I Diffusion Model. Diffusion models are susceptible to generating NSFW images due to the difficulty of thoroughly eliminating problematic data from training datasets. Recent studies have explored the potential for adversarial prompts to manipulate these models to create inappropriate images (Zhuang et al., 2023; Kou et al., 2023; Tsai et al., 2023; Ma et al., 2024; Yang et al., 2024c). For example, QF-Attack (Zhuang et al., 2023) generates adversarial prompts by minimizing the cosine distance between the features of the original prompts and those of target prompts extracted by the text encoder. Similarly, Ring-A-Bell (Tsai et al., 2023) uses steering vectors (Subramani et al., 2022) representing unsafe concepts as optimization targets for adversarial prompts. This method effectively circumvents concept removal techniques (Gandikota et al., 2023; 2024; Pham et al., 2024). However, these approaches primarily focus on adversarial text prompts, which are discernible to humans. Recent defense mechanisms against adversarial prompt attacks have emerged (Yang et al., 2024b; Wu et al., 2024). For instance, GuardT2I (Yang et al., 2024b) employs LLMs to convert encoded features of prompts back into plain texts, enabling the identification of malicious intent by distinguishing between adversarial and typical NSFW prompts.
I2I Diffusion Models. Diffusion models are employed primarily for creating new images based on textual prompts, known as T2I diffusion models (Rombach et al., 2022; Ramesh et al., 2022). More recently, researchers have discovered that these models can also modify existing images based on text instructions (Meng et al., 2021; Brooks et al., 2023; Parmar et al., 2023; Nguyen et al., 2023). SDEdit (Meng et al., 2021) changes the input from random noise to a noisy image in the inference stage, while maintaining the structure and training methodology of existing T2I models. Building on this, pix2pix-zero (Parmar et al., 2023) achieves I2I translation by preserving the input image’s cross-attention maps throughout the diffusion process. InstructPix2Pix (Brooks et al., 2023) and Visual Instruction Inversion (Nguyen et al., 2023) use images as a secondary condition alongside text, combining their features with the intermediate latent vector to enhance image editing precision. Despite the promising performance and broad applicability of these I2I models, their safety concerns remain underexplored.
3 Method
In this section, we investigate the potential safety concerns associated with diffusion models in the context of both adversarial prompt and image attacks. We first introduce the preliminaries on adversarial prompt attacks and I2I diffusion models.
3.1 Preliminaries
@@
| Sneaky | The intensity of their describing was almost too much to bear |
@@
Evaluation Using Text Filters. Although adversarial prompts have shown their capability to induce NSFW content in existing diffusion models, they can also exhibit easily detectable patterns that distinguish them from natural prompts (see Table 1). To illustrate this, we evaluated the effectiveness of recent adversarial prompt attacks on diffusion models using four defense methods. Specifically, the Perplexity Filter calculates the perplexity of the prompts using an LLM to identify adversarial prompts with abnormally high perplexity (Alon & Kamfonas, 2023). The Keyword Filter identifies NSFW prompts by detecting keywords that are in a predefined list, while the LLM Filter uses an LLM to detect both NSFW terms and non-sensical strings that may be generated by adversarial attacks. Lastly, the Embedding Filter maps input prompts into a latent space using a trained model, identifying adversarial prompts that are close to NSFW concepts but distant from safe concepts (Liu et al., 2024). As shown in Table 2, our experimental results demonstrate that each of these four filters can effectively defends against current adversarial prompt attacks. Even using the simplest text filters such as perplexity can significantly reduce the ASR of adversarial prompt attacks by around on average. We also tried the MMA-Mask attack (Yang et al., 2024a) (which is based on MMA but further removes any NSFW-related keywords) in the adversarial prompts to make the attacks more covert. The results suggest that it can only bypass the Keyword Filter, but still fails to evade the remaining three filters, particularly the LLM filter, which reduces the ASR to around .
@@
I2I Diffusion Models. I2I diffusion models for image editing take both a text prompt and an image as input. Typically, a pre-trained CLIP (Radford et al., 2021) text encoder transforms the text prompt into the text feature , while the input image is encoded into a latent feature by the encoder of a variational autoencoder (VAE) (Kingma, 2013). The diffusion process consists of timesteps, starting from random latent noise . At each timestep , a model is used to predict the noise and update the latent feature from to .
@@ 3.2 AdvI2I Framework
NSFW Concept Vector Extraction. Existing research has shown that it is possible to extract an embedding vector that represents a certain concept (Tsai et al., 2023; Ma et al., 2024) with a pair of contrastive prompts. Here we aim to extract an NSFW concept vector (e.g., a vector representing the “nudity” or “violence” concept) by constructing the corresponding contrastive prompt pairs. Specifically, the contrastive prompts consist of two sets: , which contains prompts explicitly incorporating the NSFW concept (e.g., “Let the woman naked in the car”), and , which does not contain the NSFW concept (e.g., “Let the woman in the car”). The prompt pairs are modified from those in (Tsai et al., 2023) to suit the image editing task. Then, given the text encoder , the NSFW concept can be extracted as follows:
@@
Now our goal here is to train the image generator to produce adversarial images that can lead the diffusion model to generate NSFW content while ensuring that the generated image remains visually similar to the original image. Let us denote as our generator (parameterized by ) which takes a benign image and generates an adversarial image . Unlike traditional generator training approaches (Naseer et al., 2021) that use U-Net (Ronneberger et al., 2015) or ResNet (He et al., 2016) architectures, we leverage a pre-trained VAE as the adversarial image generator to ensure greater similarity between the adversarial and original images.
@@
The constraint in Eq. (2) is to ensure that the generated image also stays close to the original image . To solve this constraint optimization problem, we apply a clipping function to the generated adversarial image, ensuring that the difference between and the input image remains within the predefined noise bound after each update step. In practice, we set in Eq. (2) since the latent feature at the final timestep111The denoising process start at timestep T and end at timestep 1. directly influences the content of the generated image.
@@
4 Experiments
4.1 Experimental Settings
Datasets. To train the adversarial noise generator and evaluate the effectiveness of AdvI2I, we construct a image-text pair dataset. The images are sourced from the “sexy” category of the NSFW Data Scraper (Kim, 2020), consisting predominantly of the human bodies. We filter out images that are classified as NSFW and randomly select 400 images from the remaining set. Additionally, 30 text prompts are generated for image editing using ChatGPT-4o (OpenAI, 2024). Then, we randomly select 200 images and 10 text prompts from each set to construct 2000 image-text pair samples, in which 1800 samples are used for training adversarial image generators and the remaining 200 samples are for evaluation.
Diffusion Models. Our experiments leverage two diffusion models. The first model, InstructPix2Pix, is modified and finetuned from SDv1.5. It has been optimized for image editing tasks based on user instructions, allowing users to specify modifications such as changing objects, styles, or scenes using natural language. The second model, SDv1.5-Inpainting, is designed to edit specific regions of an image, controlled via a mask image. We also evaluate the transferability of AdvI2I from SDv1.5-Inpainting to other SD inpainting models. The results are shown in Appendix A.1.
Baselines. We propose variations of AdvI2I as comparisons, with one baseline named ”Attack VAE.” Attack VAE modifies the loss function to generate adversarial images by only utilizing the image encoder and decoder of the diffusion model. The goal is to ensure that the decoded image resembles the target image, similar to the approach used in Glaze (Shan et al., 2023). Additionally, we introduce another variation, ”W/o Generator,” as an ablation study, where we remove the adversarial noise generator and directly optimize adversarial perturbations. For further results and analysis, please refer to Appendix A.2. In addition, we incorporate MMA-Diffusion (Yang et al., 2024a), which originally utilizes text and image modalities to generate NSFW content while evading post-hoc safety filters. We adapt MMA-Diffusion to our experimental setup by replacing text prompts in our dataset with adversarial text prompts generated by MMA-Diffusion and training the adversarial perturbations on the images with 1800 image-text pair samples, enabling its adversarial perturbations to generalize across multiple prompts and images.
Defense Strategies. We evaluate the robustness of the AdvI2I attack under three different types of defense strategies. Specifically, Safe Latent Diffusion (SLD) (Schramowski et al., 2023) and Negative Prompt (SD-NP) (Rombach et al., 2022) are popular concept removal methods applied during the inference process. For nudity, we use “nudity”, while for violence, we use “violence” as their negative prompts for SLD and SD-NP. Gaussian Noising (GN) (Hönig et al., 2024) is a pre-process defense that adds Gaussin noise to the input images. Here we use the same noise bound as the adversarial noise. Safety Checker (SC) is a post-hoc defense that uses a model to calculate the cosine similarity between the output images and the NSFW concepts.
Evaluation Metric. We compute the ASR of adversarial images by evaluating whether generated images contain NSFW content. Following (Tsai et al., 2023), to classify whether the generated images have nudity content, we employ the NudeNet (nud, 2023) detector. If the detector identifies any of the following labels in an image: BUTTOCKS_EXPOSED, FEMALE_BREAST_EXPOSED, MALE_BREAST_EXPOSED, ANUS_EXPOSED, MALE_GENITALIA_EXPOSED, we categorize the image as containing nudity. To assess whether the images contain other inappropriate content such as violence, we use the Q16 classifier (Schramowski et al., 2022).
4.2 Results and Analysis

### v3

@@ Abstract
Recent advances in diffusion models have significantly enhanced the quality of image synthesis, yet they have also introduced serious safety concerns, particularly the generation of Not Safe for Work (NSFW) content. Previous research has demonstrated that adversarial prompts can be used to generate NSFW content. However, such adversarial text prompts are often easily detectable by text-based filters, limiting their efficacy. In this paper, we expose a previously overlooked vulnerability: adversarial image attacks targeting Image-to-Image (I2I) diffusion models. We propose AdvI2I, a novel framework that manipulates input images to induce diffusion models to generate NSFW content. By optimizing a generator to craft adversarial images, AdvI2I circumvents existing defense mechanisms, such as Safe Latent Diffusion (SLD), without altering the text prompts. Furthermore, we introduce AdvI2I-Adaptive, an enhanced version that adapts to potential countermeasures and minimizes the resemblance between adversarial images and NSFW concept embeddings, making the attack even more resilient against defenses. Through extensive experiments, we demonstrate that both AdvI2I and AdvI2I-Adaptive can effectively bypass current safeguards, highlighting the urgent need for stronger security measures to address the misuse of I2I diffusion models. The code is available at https://github.com/Spinozaaa/AdvI2I.
CAUTION: This paper contains explicit content that may be disturbing to some readers.
@@ 1 Introduction
Recently, diffusion models have made significant strides in the domain of image synthesis, demonstrating their ability to produce high-quality images (Rombach et al., 2022; Zhang et al., 2023). However, these advancements have also raised significant ethical and safety concerns. Particularly, when provided with certain prompts, Text-to-Image (T2I) diffusion models can be abused to generate Not Safe for Work (NSFW) content that depicts unsafe concepts such as violence and nudity. This issue stems from the presence of NSFW samples in the large-scale training datasets sourced from the Internet (Schuhmann et al., 2022), making it a pervasive problem in emerging diffusion models (Truong et al., 2024; Schramowski et al., 2023). Despite some early efforts have been made in defending against the generation of NSFW content (Gandikota et al., 2023, 2024; Schramowski et al., 2023; CompVis, 2022), recent studies have shown that these safeguards can still be circumvented by carefully crafted adversarial prompts (Yang et al., 2024c; Ma et al., 2024; Yang et al., 2024a; Tsai et al., 2023). As a result, malicious users can exploit these models to generate NSFW images for unethical purposes.
While adversarial prompts present a notable risk to the generation safety of diffusion models, their Achilles’ heel lies in that such attacks work by changing the input text prompt, which can exhibit easily detectable patterns that distinguish them from natural prompts. Specifically, we applied four types of simple filters (perplexity filter, keyword filter, embedding filter and large language model (LLM) filter) to a range of adversarial prompt attacks (Zhuang et al., 2023; Kou et al., 2023; Tsai et al., 2023; Ma et al., 2024; Yang et al., 2024c), and found that even the simplest filters can effectively identify adversarial prompts from normal ones in most cases (see more detailed in Section 3.1). Notably, a naive perplexity filter can (on average) reduce the attack success rate (ASR) of adversarial prompts by , while LLM as the safety filter reduces the ASR to under .
This suggests that adversarial text prompts can be identified, which means that diffusion models can reject generating images with such queries. However, the new question is:
@@
In this work, we provide a negative answer to this question. We reveal the risk of adversarial images that can also induce diffusion models to generate NSFW images, which has not been well explored in previous research. We propose a framework named AdvI2I to demonstrate the effectiveness of such an attack on the Image-to-Image (I2I) diffusion model, alerting the community to adversarial attacks from not only the prompt but also the image condition side. In addition to text prompts, I2I diffusion models conventionally utilize an image as a conditioning input. By leveraging adversarial images, attackers can induce the diffusion model to generate NSFW images. For example, an image of the president can be manipulated to depict nudity. Moreover, this attack can bypass existing defense mechanisms designed for diffusion models, revealing a significant yet underexplored security vulnerability in this domain. By circumventing these defenses, AdvI2I can effectively expose the inherent risks present in I2I models, highlighting their susceptibility to generating NSFW content under adversarial influence.
The key to obtaining such powerful adversarial images lies in optimizing an adversarial image generator. The optimization target is the denoised latent feature in the diffusion process. Given that the feature is influenced by both the image and text conditions, AdvI2I transforms the NSFW concept from the text embedding space into the adversarial perturbation on images, enabling it to guide the model in generating NSFW content. Additionally, to further explore the efficacy of such adversarial attack under potential defenses, we propose a modified attack approach named AdvI2I-Adaptive. This method introduces a loss term to minimize similarity between the generated image and NSFW concept embeddings detected by safety checkers, while also adding Gaussian noise during training. By incorporating these adaptive elements, AdvI2I-Adaptive enhances the robustness of adversarial attacks against current defense measures, significantly amplifying the threat posed by adversarial images in I2I diffusion models.
Our contributions are summarized as follows.
@@
By highlighting the risk of adversarial attacks from image conditions, raising awareness within the research community about the potential dangers of such attacks on diffusion models. Our findings underscore the inherent capability of these models to generate NSFW content under adversarial influence, emphasizing the need for further research into robust defense mechanisms.
@@
| Sneaky | The intensity of their description was almost too much to bear |
@@
2 Related Work
Adversarial Attack and Defense in T2I Diffusion Models. Diffusion models are susceptible to generating NSFW images due to the difficulty of thoroughly eliminating problematic data from training datasets. Recent studies have explored the potential for adversarial prompts to manipulate these models to create inappropriate images (Zhuang et al., 2023; Kou et al., 2023; Tsai et al., 2023; Ma et al., 2024; Yang et al., 2024c). For example, QF-Attack (Zhuang et al., 2023) generates adversarial prompts by minimizing the cosine distance between the features of the original prompts and those of target prompts extracted by the text encoder. Similarly, Ring-A-Bell (Tsai et al., 2023) uses steering vectors (Subramani et al., 2022) representing unsafe concepts as optimization targets for adversarial prompts. This method effectively circumvents concept removal techniques (Gandikota et al., 2023, 2024; Pham et al., 2024). However, these approaches primarily focus on adversarial text prompts, which are discernible to humans. Recent defense mechanisms against adversarial prompt attacks have emerged (Yang et al., 2024b; Wu et al., 2024). For instance, GuardT2I (Yang et al., 2024b) employs LLMs to convert encoded features of prompts back into plain texts, enabling the identification of malicious intent by distinguishing between adversarial and typical NSFW prompts.
I2I Diffusion Models. Diffusion models are employed primarily for creating new images based on textual prompts, known as T2I diffusion models (Rombach et al., 2022; Ramesh et al., 2022). More recently, researchers have discovered that these models can also modify existing images based on text instructions (Meng et al., 2021; Brooks et al., 2023; Parmar et al., 2023; Nguyen et al., 2023). SDEdit (Meng et al., 2021) changes the input from random noise to a noisy image in the inference stage, while maintaining the structure and training methodology of T2I models. Building on this, pix2pix-zero (Parmar et al., 2023) achieves I2I translation by preserving the input image’s cross-attention maps throughout the diffusion process. InstructPix2Pix (Brooks et al., 2023) and Visual Instruction Inversion (Nguyen et al., 2023) use images as a secondary condition alongside text, combining their features with the intermediate latent vector to enhance image editing precision. Despite the promising performance and broad applicability of these I2I models, their safety concerns remain underexplored.
3 Method
In this section, we investigate the potential safety concerns associated with diffusion models in the context of both adversarial prompt and image attacks. We first introduce the preliminary experiments on adversarial prompt attacks and the structure of I2I diffusion models.
3.1 Preliminaries
@@
Evaluation Using Text Filters. Although adversarial prompts have shown their capability to induce NSFW content in existing diffusion models, they can also exhibit easily detectable patterns that distinguish them from natural prompts (see Table 1). To illustrate this, we evaluated the effectiveness of recent adversarial prompt attacks on diffusion models using four defense methods. Specifically, the Perplexity Filter calculates the perplexity of the prompts using an LLM to identify adversarial prompts with abnormally high perplexity (Alon & Kamfonas, 2023). The Keyword Filter identifies NSFW prompts by detecting keywords that are in a predefined list, while the LLM Filter uses an LLM to detect both NSFW terms and non-sensical strings that may be generated by adversarial attacks. Lastly, the Embedding Filter maps input prompts into a latent space using a trained model, identifying adversarial prompts that are close to NSFW concepts but distant from safe concepts (Liu et al., 2024). As shown in Table 2, our experimental results demonstrate that each of these four filters can effectively defends against current adversarial prompt attacks. Even using the simplest text filters such as perplexity can significantly reduce the ASR of adversarial prompt attacks by around on average. We also tried the MMA-Mask attack (which is based on MMA (Yang et al., 2024a) but further removes any NSFW-related keywords) in the adversarial prompts to make the attacks more covert. The results suggest that it can only bypass the Keyword Filter, but still fails to evade the remaining three filters, particularly the LLM filter, which reduces the ASR to around .
@@
I2I Diffusion Models. I2I diffusion models for image editing take both a text prompt and an image as inputs. Typically, a pre-trained CLIP (Radford et al., 2021) text encoder transforms the text prompt into the feature vector , while the input image is encoded into a latent feature by the encoder of a variational autoencoder (VAE) (Kingma, 2013). Then, the diffusion process is applied, which consists of timesteps, starting from a random latent noise . At each timestep , a model is used to predict the noise and update the latent feature from to .
@@ 3.2 AdvI2I Framework
NSFW Concept Vector Extraction. Existing research has shown that it is possible to extract an embedding vector that represents a certain concept (Tsai et al., 2023; Ma et al., 2024) with a pair of contrastive prompts. Here we aim to extract an NSFW concept vector (e.g., an intermediate feature vector representing the “nudity” or “violence” concept) by constructing the corresponding contrastive prompt pairs. Specifically, the contrastive prompts consist of two sets: , which contains prompts explicitly incorporating the NSFW concept (e.g., “Let the woman naked in the car”), and , which does not contain the NSFW concept (e.g., “Let the woman in the car”). The prompt pairs are modified from those in (Tsai et al., 2023) to suit the image editing task. Then, given the text encoder , the NSFW concept can be extracted as follows:
@@
Then our goal is to train the image generator to produce adversarial images that can lead the diffusion model to generate NSFW content while ensuring that the generated image remains visually similar to the original image. Let us denote as our generator (parameterized by ), which takes a benign image and generates an adversarial one . Unlike traditional adversarial image generators on the classification task (Naseer et al., 2021) that use U-Net (Ronneberger et al., 2015) or ResNet (He et al., 2016) models, we leverage a pre-trained VAE to ensure greater similarity between the adversarial and original images.
@@
The constraint in Eq. (2) is to ensure that the generated image also stays close to the original image . To solve this constraint optimization problem, we apply a clipping function to the generated adversarial image, ensuring that the difference between and the input image remains within the predefined noise bound after each update step. In practice, we set in Eq. (2) since the latent feature at the final timestep111The denoising process starts at timestep and ends at . directly influences the content of the generated image.
@@
4 Experiments
4.1 Experimental Settings
Datasets. To train the adversarial noise generator and evaluate the effectiveness of AdvI2I, we construct an image-text dataset (i.e., one sample includes an image and a text prompt). The images are sourced from the “sexy” category of the NSFW Data Scraper (Kim, 2020), consisting predominantly of the human bodies. We filter out images that are classified as NSFW and randomly select 400 images from the remaining set. Additionally, 30 text prompts are generated for image editing using ChatGPT-4o (OpenAI, 2024). Then, we randomly select 200 images and 10 text prompts from each set to construct 2000 image-text samples, in which 1800 samples are used for training adversarial image generators and the remaining 200 samples are for evaluation.
Diffusion Models. Our experiments leverage two diffusion models. The first model, InstructPix2Pix, is modified and finetuned from SDv1.5. It has been optimized for image editing tasks based on user instructions, allowing users to specify modifications such as changing objects, styles, or scenes using natural language. The second model, SDv1.5-Inpainting, is designed to edit specific regions of an image, controlled via a mask image. We also evaluate the transferability of AdvI2I from SDv1.5-Inpainting to other SD inpainting models. The results are shown in Appendix B.
Baselines. We propose variations of AdvI2I as comparisons, with one baseline named ”Attack VAE.” Attack VAE modifies the loss function to generate adversarial images by only utilizing the image encoder and decoder of the diffusion model. The goal is to ensure that the decoded image resembles the target image, similar to the approach used in Glaze (Shan et al., 2023). Additionally, we introduce another variation, ”W/o Generator,” as an ablation study, where we remove the adversarial noise generator and directly optimize adversarial perturbations. For further results and analysis, please refer to Appendix C. In addition, we incorporate MMA-Diffusion (Yang et al., 2024a), which originally utilizes text and image modalities to generate NSFW content while evading post-hoc safety filters. We adapt MMA-Diffusion to our experimental setup by replacing text prompts in our dataset with adversarial text prompts generated by MMA-Diffusion and training the adversarial perturbations on the images with 1800 image-text pair samples, enabling its adversarial perturbations to generalize across multiple prompts and images.
Defense Strategies. We evaluate the robustness of the AdvI2I attack under three different types of defense strategies. Specifically, Safe Latent Diffusion (SLD) (Schramowski et al., 2023) and Negative Prompt (SD-NP) (Rombach et al., 2022) are popular concept removal methods applied during the inference process. We modify the parameters of the SLD to adapt to I2I diffusion models. Please refer to Appendix A for more details. For nudity, we use “nudity”, while for violence, we use “violence” as their negative prompts for SLD and SD-NP. Gaussian Noising (GN) (Hönig et al., 2024) is a pre-process defense that adds Gaussin noise to the input images. Here we use the same noise bound as the adversarial noise. Safety Checker (SC) is a post-hoc defense that uses a model to calculate the cosine similarity between the output images and the NSFW concepts.
Evaluation Metric. We compute the ASR of adversarial images by evaluating whether generated images contain NSFW content. Following (Tsai et al., 2023), to classify whether the generated images have nudity content, we employ the NudeNet (nud, 2023) detector. If the detector identifies any of the following labels in an image: BUTTOCKS_EXPOSED, MALE_GENITALIA_EXPOSED, FEMALE_BREAST_EXPOSED, ANUS_EXPOSED, MALE_BREAST_EXPOSED, we categorize the image as containing nudity. To assess whether the images contain other inappropriate content such as violence, we use the Q16 classifier (Schramowski et al., 2022).
4.2 Results and Analysis